OpenStudy (anonymous):

HALP!! v= the square root of 2as Solve for s (I said the square root of because idk how to put that symbol)

OpenStudy (anonymous):

\[v=\sqrt{2as}\]

OpenStudy (anonymous):

\[a=\frac{ 2 }{ s }\]

OpenStudy (anonymous):

Please don't give out direct answers :) @SurpriseBreeze Instead, try engaging the asker so they can learn better and understand where your answer is coming from. Not only is it against the CoC, but also prohibiting learning. As the saying goes… \(\\\color{darkblue}{Give~a~man~a~fish,~he~eats~for~a~day.}\) \(\\\color{pink}{Teach~a~man~to~fish,~he~eats~for~a~lifetime.}\)

OpenStudy (anonymous):

Type it into a calculator.. :)

OpenStudy (anonymous):

Thank you however I have to show my work so if you could show me the steps that be amazing

OpenStudy (anonymous):

I just used mathway :c They charge you to see the steps..

OpenStudy (anonymous):

yeah I use mathway too

OpenStudy (anonymous):

oh well thanks for trying!

OpenStudy (anonymous):

next person pretty please

geerky42 (geerky42):

Then mathway sucks. Where is variable v anyway?

OpenStudy (anonymous):

v is on the other side I have to somehow end up with s in its place

OpenStudy (anonymous):

SO it has to be s=blah blah

geerky42 (geerky42):

I know, I am talking about SurpriseBreeze's answer. Ok your goal is to isolate s. \(v = \sqrt{2as}\) We need to get rid of square root. Any idea what we need to do to both sides?

OpenStudy (anonymous):

Square it

geerky42 (geerky42):

Right. \(v^2 = (\sqrt{2as})^2\) \(v^2 = 2as\) Now what? How to move \(2a\) to other side?

OpenStudy (anonymous):

divide

geerky42 (geerky42):

Yeah. \(\large \boxed{s = \dfrac{v^2}{2a}}\) That it.

OpenStudy (anonymous):

Ok thank you so much!!!

OpenStudy (anonymous):

Tiger-Algebra shows the work… wolfram alpha is better though but you do have to pay when getting steps.